A sequence (In) is defined by Io = 1 and for every n E N* , In = Integral from 0→1 [dx / ( 1 + x^2 )^n ]
(a) Justify that In is welled defined and determine its sign.
(b) Show that In is a decreasing sequence.
1
Expert's answer
2021-02-02T05:27:53-0500
  1. As "\forall n \in \mathbb{N}, \frac{1}{(x^2+1)^n}" is a continuous function on "[0;1]" and "[0;1]" is a compact interval, "\forall n\in \mathbb{N}, I_n" is well defined. In addition, as "\frac{1}{(1+x^2)^n} \geq 0", the integral "I_n \geq 0".
  2. As "\frac{1}{1+x^2} \leq 1", "\frac{1}{(1+x^2)^{n+1}} \leq \frac{1}{(1+x^2)^n}" and thus "I_{n+1} \leq I_n", the sequence is non-increasing. To show that it is strictly decreasing, it is enough to decompose this integral into two parts : "I_n = \int_0^{1/2}\frac{1}{(x^2+1)^n} dx + \int_{1/2}^1 \frac{1}{(x^2+1)^n} dx", the first part is non-increasing due to the same inequality as before, in the second part "\frac{1}{x^2+1} \leq \frac{4}{5} < 1" and thus "\frac{1}{(1+x^2)^{n+1}} < \frac{1}{(1+x^2)^n}, x\in[1/2;1]" and so the second part is strictly decreasing. The sum of a non-increasing and a strictly secreasing sequences is a strictly decreasing sequence and thus "I_n" is decreasing.